OpenStudy (anonymous):

Solve for variable 3w=4 1/5

OpenStudy (kewlgeek555):

Okay. This is the equation, right?:\[3w=4\frac{ 1 }{ 5 }\]

OpenStudy (anonymous):

yup

OpenStudy (kewlgeek555):

e.e

OpenStudy (kewlgeek555):

Okay. So it has a simple start. What equals four and one half? Put it in decimal form. This is easy. Do you know how to do this @reyna78724 ?

OpenStudy (nirmalnema):

3w=21/5 w=21/15 w=7/5

OpenStudy (anonymous):

Yeah that's it

OpenStudy (kewlgeek555):

Hm...it seems that @nirmalnema got a slightly different answer than I did. I might have done something wrong or he/she might have done something wrong.

OpenStudy (kewlgeek555):

I don't know. I know that I know that 4 and one half is 4.5. That made the equation 3w = 4.5. Now you have to divide. 4.5 divided by 3 is 1.5. Now, @nirmalnema got 1.4...Hmm.

OpenStudy (nirmalnema):

@kewlgeek555 21/ 5 is = 4.2 not 4.5

OpenStudy (kewlgeek555):

21? Where did you get 21 from?

OpenStudy (nirmalnema):

by solving 4 1/5 4*5+1

OpenStudy (anonymous):

Well the answer chose,s are -7/5 7/5,63/5 -63/5

OpenStudy (nirmalnema):

so 2nd is correct

OpenStudy (anonymous):

Okay what about this is -m/4.2=5

OpenStudy (nirmalnema):

m=-21

OpenStudy (anonymous):

And one more g/24=-2

OpenStudy (nirmalnema):

g=-48

OpenStudy (nirmalnema):

@reyna78724 is that correct